OpenStudy (anonymous):

Use change of base to rewrite the expressio: log7 25 I wrote: 7^x = 25 Is this correct? Second question: Simplify x^2-2x-15/(x-6)(x-3) * x^2-6/(x-5) Any ideas or help?

OpenStudy (anonymous):

change of base \[\log_7(25)=\frac{\ln(25)}{\ln(7)}\] what you did was write it in equivalent exponential form

OpenStudy (anonymous):

So what you wrote exactly would be the right answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

question asks for change of base, but they didn't specify the base ( or if they did, i can't see it)

OpenStudy (anonymous):

the first one, you are right. second: